OpenStudy (anonymous):

HELP WITH SLOPE.The equation of line XY is (y-3) = -2/3 (x - 4). What is the slope of a line perpendicular to XY? 3/2?

OpenStudy (anonymous):

@sammixboo

OpenStudy (anonymous):

@nincompoop

OpenStudy (godlovesme):

what do u think the answer is @Tatsurion

OpenStudy (anonymous):

I think it's 3/2

OpenStudy (godlovesme):

u r right *medals*

OpenStudy (anonymous):

yay :D

OpenStudy (godlovesme):

awesome job and thx for the medal :)))))

OpenStudy (anonymous):

The equation of line EF is y = 1/2 x + 6. Write an equation of a line parallel to line EF in slope-intercept form that contains point (0, -2). What about this? y=1/2x-2?

OpenStudy (anonymous):

:) I like giving medals In history I won't stop helping until i am The Immortal One >:D

OpenStudy (godlovesme):

y-y1=m(x-x1) m is slope y1 is -2 x1 is 0 can u plug this info to the eqn^^

OpenStudy (anonymous):

Is it y = 1/2x-2?

OpenStudy (godlovesme):

@Tatsurion xDDDD the immortal one lolol :)))

OpenStudy (anonymous):

Is it y = 1/2x-2?

OpenStudy (godlovesme):

omg u r AWESOME fr!!! u don't even need help :)) wish i could medal u again >.<

OpenStudy (godlovesme):

btw u r correct :)

OpenStudy (godlovesme):

@Tatsurion